OpenStudy (anonymous):

Choose the equation below whose axis of symmetry is x = 2. y = x^2 + 4x + 2 y = x^2 − 4 y = x^2 − 2 y = x^2 − 4x + 2

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (imstuck):

I can help you unless you want to wait for jim_thompson5910

OpenStudy (anonymous):

ima say A

jimthompson5910 (jim_thompson5910):

you can go ahead IMStuck

jimthompson5910 (jim_thompson5910):

hint: the axis of symmetry of y = ax^2 + bx + c is x = -b/(2a)

OpenStudy (anonymous):

@IMStuck I would appreciate your assistance.

OpenStudy (anonymous):

I put A but I wanted to double check my answers

OpenStudy (anonymous):

sorry D the formula is -b/a that would make the -4 positive

OpenStudy (anonymous):

if you go with A (my first answer) the answer would turn negative i forgot about the formula

OpenStudy (anonymous):

Functions 1 and 2 are shown below. Function 1: f(x) = −3x2 + 2 A graph of a parabola with x intercepts of negative 0.5, 0 and 2, 0 and a vertex of 0.5, 4 is shown. Which function has a larger maximum? Type your answer as 1 or 2.

OpenStudy (anonymous):

The 2 in function 1 is an exponent btw

OpenStudy (anonymous):

-3x^2+2